(5x+1)/6=(4x-1)/2


Found 2 solutions by mananth, CenturySimons:
Answer by mananth(16949)

(5x+1)/6=(4x-1)/2
..
2*(5x+1)=6*(4x-1)
10x+2=24x-6
24x-10x=6+2
14x= 8
x= 8/14
x=4/7

Answer by CenturySimons(1)

Step 1: Multiply both sides of the equation by 6 which is the least common denominator of both fractions. This will clear the equations of the fractions.
(5x+1)/6=(4x-1)/2
6((5x+1)/6)=6((4x-1)/2)
(5x+1)=3(4x-1)
Step 2: Distribute the factor of 3 on the right to get
5x+1=12x-3
Step 2: Subtract 5x from and add 3 to both sides of the equation to get
4=7x
Step 3: Divide both sides of the equation by 7 to get
4/7= x (SOUTION)